Two angles atre complementary. The sum of the measure of the first angle and one-fourth the second angle is 69 degrees. Find the measure of the angles.
Answer by nerdybill(5404)

Two angles atre complementary. The sum of the measure of the first angle and one-fourth the second angle is 69 degrees. Find the measure of the angles.
.
Two angles are complementary if their sum equals 180 degrees.
.
Let x = measure of second angle
then
180-x = measure of first angle
.
(180-x) + (1/4)x = 69
720 - 4x + x = 276
720 - 3x = 276
-3x = -444
x = 148 degrees (second angle)
.
First angle:
180-x = 180-148 = 32 degrees (first angle)